# DECO: Life-Cycle Management of Enterprise-Grade Chatbots

### v1

@@ Abstract.
Software engineers frequently grapple with the challenge of accessing disparate documentation and telemetry data, including Troubleshooting Guides (TSGs), incident reports, code repositories, and various internal tools developed by multiple stakeholders. While on-call duties are inevitable, incident resolution becomes even more daunting due to the obscurity of legacy sources and the pressures of strict time constraints. To enhance the efficiency of on-call engineers (OCEs) and streamline their daily workflows, we introduced DECO—a comprehensive framework for developing, deploying, and managing enterprise-grade chatbots tailored to improve productivity in engineering routines. This paper details the design and implementation of the DECO framework, emphasizing its innovative NL2SearchQuery functionality and a hierarchical planner. These features support efficient and customized retrieval-augmented-generation (RAG) algorithms that not only extract relevant information from diverse sources but also select the most pertinent toolkits in response to user queries. This enables the addressing of complex technical questions and provides seamless, automated access to internal resources. Additionally, DECO incorporates a robust mechanism for converting unstructured incident logs into user-friendly, structured guides, effectively bridging the documentation gap. Feedback from users underscores DECO’s pivotal role in simplifying complex engineering tasks, accelerating incident resolution, and bolstering organizational productivity. Since its launch in September 2023, DECO has demonstrated its effectiveness through extensive engagement, with tens of thousands of interactions from hundreds of active users across multiple organizations within the company.
@@ 1. Introduction
Introduction to DECO—a framework for developing, deploying, and managing chatbots
In this paper, we introduce DECO, a framework for developing, deploying, and managing copilots chatbots designed to support the continuous evolution of a software development team’s knowledge base while simplifying access to various internal tools. Chatbots configured with DECO streamline daily tasks, improve incident management, and provide engineers with a unified access point to a wide range of internal resources and tools. DECO focuses on four key areas: (1) a Generalized Development Platform, enabling contributors to easily create and deploy new chatbots; (2) Skills Integration, allowing teams to add new capabilities that incorporate additional knowledge sources or expose skills for accessing internal tools through a chat interface; (3) Fine-Tuned Retrieval Algorithms, designed for efficiency and accuracy in production environments; and (4) Streamlined End-to-End Deployment and Lifecycle Management, offering a self-hosting model with fully automated deployment pipelines, continuous monitoring, and testing capabilities to support debugging and ongoing improvements. Additionally, DECO supports an extensible development platform that enables integration with internal tools—such as debugging utilities and monitoring systems—frequently used in the daily tasks of engineers.
@@
Using DECO, teams across various divisions at Microsoft have successfully deployed chatbots tailored to their specific roles without needing extensive expertise in large language models (LLMs) or the development of LLM applications. Within the Azure Data organization, which includes services like Azure SQL DB, Microsoft Fabric, and Microsoft Synapse, we have recorded tens of thousands of messages from hundreds of users (see Figure 1) since its launch in September 2023. From the several hundred feedback entries received (rated in stars), the average rating is 3.6. Interviews with on-call engineers (OCEs) reveal that DECO not only supports routine tasks but also significantly reduces the time required for incident triaging—by approximately 10–20 minutes per incident. This efficiency translates into annual cost savings amounting to tens of millions in the cost of goods sold (COGS) for the organization. Engineers report that these chatbots have become their primary tool for information retrieval tasks.
Challenge
DECO addresses several fundamental challenges:
@@
Generalized Framework. Traditional development of chatbot or copilot applications typically involves iterative adjustments and prompt tuning through trial and error to attain satisfactory performance. In contrast, DECO introduces an “autonomous mode” that facilitates self-onboarding for tenants, adopting a “one-size-fits-all” approach posing a higher level of development sophistication.
This generalized solution was achieved through a three-step process: (1) Pilot Testing: Initially, we crafted an end-to-end solution for a select group of pilot tenants (such as Azure SQL DB and Fabric Data Warehouse), during which we refined prompts and retrieval algorithms. This phase led to the development of a set of reusable templates. We differentiated between modules needing customization for team-specific functionalities (e.g., product-specific prompts) and those that could be standardized across tenants. (2) Generalization: We then expanded the tool to accommodate all teams (tenants) by developing it into an SDK, enabling straightforward adoption across Microsoft. (3) Customization: Anchored on a flexible base, the framework permits teams to augment the chatbot with new skills or to integrate existing skills from our library, promoting a crowdsourced development model. To boost the chatbot’s performance and contextual relevance to a specific service, tenants are required to configure a few files, such as the team name, Git repositories, and a detailed system description of their services (refer to the example in Figure 2 for Fabric Data Warehouse). The chatbot is continually enhanced through user feedback (Anonymous, 2024b) and an evolving evaluation toolkit tailored to the specific usage data of each team.
fabric_dw.md
  # Introduction
    Microsoft Fabric DW is a data analytics solution ...
  # Control Plane ...
Figure 2. Team description example
Efficiency. We propose a self-hosting model where each tenant, often with limited resources or expertise, assumes full responsibility for the end-to-end setup and maintenance, guided by our comprehensive onboarding documentation and fully automated deployment pipeline through ARM or Bicep  (Microsoft, 2024c). Our approach prioritizes simplicity, minimizing the need for extensive resource allocation and adjustments, thereby facilitating more efficient deployment and reducing maintenance costs. We employ mature, established tools that offer robust monitoring and debugging capabilities in production environments.
Knowledge Gap. DECO addresses the challenge of transforming raw incident logs into comprehensive and curated reports, thereby mitigating deficiencies in existing documentation. Recognizing the critical role that well-structured documentation plays in providing accurate context to large language models (LLMs), DECO integrates a robust mechanism for converting unstructured incident logs into user-friendly, structured guides. By automating the creation of enriched documentation from historical incident logs through scheduled preprocessing pipelines, DECO effectively bridges documentation gaps in a timely manner.
@@
Response Quality. To enhance the quality of the chatbot’s responses, we implemented several key improvements:
@@
DECO employs a hierarchical skill selection framework and integrates advanced retrieval algorithms to optimize both accuracy and functionality. The most common issues with the chatbot are not due to hallucinations—likely avoided through well-constructed prompts—but rather stem from the incorrect selection of skills or retrieval of documents. This issue is particularly prevalent when managing a large inventory of skills and documentation segments. Moreover, the documentation sources for DECO chatbots often consist of highly technical documents that appear “similar”, which complicates the retrieval process.
To manage diverse functionalities and improve response accuracy, DECO utilizes a hierarchical planner that assigns tasks to specialized sub-planners. This system categorizes skills by functionality and usage frequency to optimize deployment. Additionally, by leveraging state-of-the-art algorithms from Azure AI Search (Microsoft, 2024b), DECO significantly enhances its document retrieval capabilities from both documentation sites, Incident Management (IcM) databases, and code repositories through an innovative “NL2SearchQuery” method. This ensures that the LLM is provided with the most accurate and relevant information.
@@
DECO incorporates a comprehensive evaluation system alongside telemetry data collection pipelines to facilitate online monitoring and debugging. After deployment across multiple tenants, DECO continuously refines response quality through a robust evaluation framework. This system measures chatbot performance using various metrics, including answer relevance as evaluated by LLMs and user feedback. In addition, several specialized evaluators have been developed to verify the accuracy of individual modules within the chatbot workflow. Hosted on Azure, DECO features an extensive telemetry data collection infrastructure that not only monitors chatbot response quality but also supports an efficient logging system for prompt issue triaging.
@@
In summary, DECO is a framework designed for the self-hosting and development of production-ready chatbots. It offers substantial support to engineers in their daily and on-call tasks by efficiently surfacing relevant information, streamlining access to various internal tools, and bridging gaps between telemetry data, product documentation, and code. The main contributions of DECO are as follows:
  • An end-to-end pipeline for the development, deployment, and management of enterprise-grade chatbots that support software engineers’ daily routines and incident resolution.
@@
  • Efficient retrieval and skill selection methods for document, internal toolkit, and code access, enhancing chat quality.
  • A robust evaluation and monitoring system to facilitate continuous improvement and effective error triaging.
The remainder of the paper is organized as follows: Section 2 covers background information and data sources. Section 3 outlines the overall system design. Section 4 describes the data preprocessing pipelines for knowledge base preparation. Section 5 details the backend infrastructure, and Section 6 explains the advanced algorithms used to enhance response accuracy. Section 7 presents the frontend components. Section 8 describes the evaluation framework and experimental results. Section 9 discusses related work, and Section 10 concludes the paper.
@@ 2. Background
The primary objective of DECO is to provide a framework for setting up chatbots that offer engineers a time-efficient, interactive experience for managing daily tasks and on-call responsibilities. DECO effectively bridges gaps between product telemetry, documentation, on-call duties, code development resources, and team-specific toolkits. However, it is important to note that these chatbots are not intended to fully replace existing automated diagnostic platforms, such as decision-tree-based systems developed for specific services (Xuanhe Zhou, 2023; Zhou et al., 2023); rather, they are designed to complement them.
@@ 2.1. Data Sources
By default, DECO incorporates resources from (1) internal documentation repositories, such as troubleshooting guides; (2) historical incident databases; (3) code repositories; and (4) internal Stack Overflow. Tenants can also add additional skills to integrate more sources as needed.
@@
Historical mitigated incidents from the Incident Management (IcM) system document unexpected disruptions or outages of services or products (Chen et al., 2020). Microsoft has a sophisticated incident management system designed to support issue identification, team assignment, and reporting to help avert or mitigate future incidents  (Chen et al., 2020). Incident data, including information such as title, summary, source, and owning team, is meticulously stored in a dedicated database. This database also captures all communications between engineers, including investigation notes. These records serve not only as a comprehensive documentation resource for incident triaging but also provide invaluable guidance for broader engineering practices.
@@
Similar to IcM databases, Stack Overflow serves as an invaluable resource for Microsoft engineers, offering a wealth of community-driven knowledge and problem-solving techniques. Discussions from these forums are extracted and summarized into reports for future reference.
@@ 3. Architecture
This section provides a comprehensive overview of DECO’s architecture and design decisions. DECO comprises 4 primary components: Data Preprocessing, Backend Services, Frontend Services, and Evaluation (see Figure 3):
Figure 3. DECO architecture
Data Preprocessing: The data preprocessing pipelines are orchestrated as Azure Machine Learning experiments (Microsoft, 2024d), executed at user-defined intervals to account for information updates frequently. They draw inputs from various data sources, such as Git, Wikis, and Kusto databases, to generate summarized documents stored in Azure Blob Storage for subsequent retrieval. Multiple Azure AI Search indexes (Microsoft, 2024b) are created to reference the corresponding document blobs, enabling easy access via API calls from downstream backend modules. Table 1 compares the total indexing time of Azure AI Search against alternative vector databases for SQL documents from six repositories containing approximately 4,000 markdown files, demonstrating Azure AI Search’s high efficiency. Section 4 provides a more detailed discussion of this module.
Table 1. Indexing execution time for SQL from 6 repositories.
Vector DB Indexing Execution Time
Faiss $\sim$7 mins
ChromaDB $\sim$24 mins
ACS $\sim$6 mins
Backend Services: The backend orchestrates the Planner-Skill chat flow using PromptFlow (Microsoft, 2024n). PromptFlow simplifies the integration of LLMs, prompts, and Python code into an executable workflow structured as a Directed Acyclic Graph (DAG), coordinating input and output exchanges across various modules (e.g., LLM calls, database connections, and other APIs). Integrated closely with the Azure Machine Learning infrastructure (Microsoft, 2024d), PromptFlow enables straightforward deployment, API and flow monitoring, and greatly reduces maintenance complexity for developers. The flow is deployed as a managed online endpoint (Microsoft, 2024f) with a stateless REST API in Azure Machine Learning, optimized for real-time, high-volume, low-latency inference requests, and equipped with extensive logging and debugging capabilities. Additionally, various retrieval nodes (authored as context-retrieval skills) can be configured alongside other skills to automatically invoke internal toolkits based on the planner’s node output. Sections 5 and 6 provide further details on these modules.
Frontend Services: The frontend manages user authentication and authorization through Azure Active Directory (Azure Intra ID) and handles user interactions. It manages session memory, building the full chat history and sending it to the backend’s (stateless) REST API to enable a continuous conversation experience. Session state and extracted memory data are stored in Azure Blob Storage, while Azure Log Analytics automatically stores telemetry data for easy monitoring and review. This data is later used for feedback learning (Anonymous, 2024b) and evaluation (see Section 8). A web application is deployed to interact with the Azure Bot Framework, supporting two main user interfaces: a web-based platform and a Microsoft Teams integration, including group chat and session-sharing functionalities.
Implementing a stateless REST API in the backend enables the RAG/skill/agent functions to operate independently of session state management and authentication, offering two key advantages:
@@
Evaluation: A comprehensive assessment of the entire system is conducted to monitor and evaluate real user interactions as well as individual modules, ensuring a high-quality chat experience. Section 8 discusses this module in more detail.
Figure 4. Preprocessing pipelines for IcM
@@ 4. Data Preprocessing
In this section, we describe the preprocessing pipeline for three data sources as examples: IcM (historical mitigated incidents), documentation, and code repositories. These pipelines are designed to support context-retrieval skills in the backend, such as get_tsg, get_icm and get_code. Additional pipelines can be created using the provided template, such as the processing pipeline to integrate information from sources like Stack Overflow or internal email exchanges. In the backend, AI search indexes created through these pipelines can be accessed by additional context-retrieval skills. These regularly scheduled asynchronous preprocessing jobs enable the automatic ingestion and indexing of data to account for any updates promptly (e.g., every 6 hours).

### v2

# DECO: Life-Cycle Management of Production-Scale Copilots [Industry]

@@ Abstract.
Software engineers frequently grapple with the challenge of accessing disparate documentation and telemetry data, including TroubleShooting Guides (TSGs), incident reports, code repositories, and various internal tools developed by multiple stakeholders. To enhance the efficiency of on-call engineers (OCEs) and streamline their daily workflows, we introduced DECO—a comprehensive framework for developing, deploying, and managing copilots tailored to improve productivity in engineering routines. This paper details the design and implementation of the DECO framework, emphasizing its innovative NL2SearchQuery functionality and a lightweight agentic framework. These features support efficient and customized retrieval-augmented-generation (RAG) algorithms that not only extract relevant information from diverse sources but also select the most pertinent skills in response to user queries. This enables the addressing of complex technical questions and provides seamless, automated access to internal resources. Additionally, DECO incorporates a robust mechanism for converting unstructured incident logs into user-friendly, structured guides, effectively bridging the documentation gap. Since its launch in September 2023, DECO has demonstrated its effectiveness through widespread adoption, enabling tens of thousands of interactions and engaging hundreds of monthly active users (MAU).
@@
PVLDB Reference Format:
Yiwen Zhu, Mathieu Demarne, Kai Deng, Wenjing Wang, Nutan Sahoo, Divya Vermareddy, Hannah Lerner, Yunlei Lu, Swati Bararia, Anjali Bhavan, William Zhang, Xia Li, Katherine Lin, Miso Cilimdzic, Subru Krishnan. PVLDB, 14(1): XXX-XXX, 2020.
doi:XX.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 14, No. 1 ISSN 2150-8097.
doi:XX.XX/XXX.XX
@@ 1. Introduction
Introduction to DECO—a framework for developing, deploying, and managing copilots
In this paper, we introduce DECO, a framework for developing, deploying, and managing copilot applications designed to support the continuous evolution of a software development team’s knowledge base while simplifying access to various internal tools. Copilots configured with DECO streamline daily tasks, improve incident management, and provide engineers with a unified access point to a wide range of internal resources and tools. DECO focuses on four key areas: (1) a Generalized Development Platform, enabling contributors to easily create and deploy new copilots; (2) Skills/Agent Integration, allowing teams to add new capabilities that incorporate additional knowledge sources or expose skills for accessing internal tools through a chat interface; (3) Fine-Tuned Retrieval Algorithms, designed for efficiency and accuracy in production environments; and (4) Streamlined End-to-End Deployment and Lifecycle Management, offering a self-hosting model with fully automated deployment pipelines, continuous monitoring, and testing capabilities to support debugging and ongoing improvements. Additionally, DECO supports an extensible development platform that enables integration with internal tools—such as debugging utilities and monitoring systems—frequently used in the daily tasks of engineers.
@@
Using DECO, dozens of teams across various divisions at Microsoft have successfully deployed copilots tailored to their specific roles. Within the Azure Data organization—which includes services such as Azure SQL DB, Microsoft Fabric, and Microsoft Synapse—we have recorded tens of thousands of messages and over 500 monthly active users (MAU) [^1] Many users are on-call engineers, typically working in rotation. (see Figure 1). Across approximately 1,500 monthly sessions, conversations average 2.2 rounds, indicating typical multi-turn interactions. From the several hundred feedback entries received (rated in stars), the average rating is 3.6. Interviews with on-call engineers (OCEs) reveal that DECO not only supports routine tasks but also significantly reduces the time required for incident triaging—by approximately 10–20 minutes per incident.
Introduction to DECO
DECO introduces the following innovations:
@@
Generalized Framework. Traditional development of copilot applications typically involves iterative adjustments and prompt tuning through trial and error to attain satisfactory performance. In contrast, DECO introduces an “autonomous mode” that facilitates self-onboarding for tenants, adopting a “one-size-fits-all” approach posing a higher level of development sophistication.
Knowledge Gap. Recognizing the challenges of maintaining high-quality documentation and its critical role in providing accurate context to large language models (LLMs), DECO implements a robust mechanism to transform large volumes of incident logs into user-friendly, structured guides, enriching the knowledge base.
@@
Response Quality. To enhance the quality of the copilot’s responses, we implemented several key improvements:
@@
DECO employs an agentic framework for skill and agent selection, integrating advanced retrieval algorithms to optimize both accuracy and functionality. The most common issues with the copilot are not due to hallucinations (see Section 8.2, where only 0.1% of responses are identified as having grounding issues), but rather document relevancy problems. This challenge becomes particularly pronounced when managing a large inventory of skills and documentation segments. Furthermore, the documentation sources for DECO copilots often consist of highly technical materials with similar content, which further complicates the retrieval process.
To manage diverse functionalities and enhance response accuracy, DECO employs an agentic framework to select the appropriate agent and corresponding skills for addressing user queries. Compared to using a single planner to choose from a larger set of skills, this sequential selection approach significantly improves skill selection accuracy. Additionally, we develop an an innovative “NL2SearchQuery” method to enhance document retrieval from documentation sites, Incident Management (IcM) databases, and other text-based sources. For code search, we introduce the Language-Agnostic RA (LA-RAG), which preprocesses large code repositories containing multiple languages while preserving code component relationships. The solutions are not only lightweight (requiring no fine-tuning or post-training) but are also supported by existing commercial indexing services, such as Azure AI Search (Microsoft, 2024b).
@@
DECO incorporates a comprehensive evaluation system alongside telemetry data collection pipelines to facilitate online monitoring and debugging. Following deployment across multiple tenants, we continuously monitor response quality through a robust evaluation framework. This system assesses copilot performance using diverse metrics, including answer relevance as judged by LLMs and user feedback. Additionally, specialized evaluators are implemented to verify the accuracy of individual modules within the copilot workflow and identify opportunities for improvement. Hosted on Azure, DECO features a comprehensive telemetry infrastructure that not only tracks copilot response quality but also supports efficient logging for rapid issue triaging.
@@
The main contributions of DECO are as follows:
  • An end-to-end pipeline for the development, deployment, and management of production-grade copilots that support software engineers’ daily routines and incident resolution.
@@
  • An efficient retrieval algorithm utilizing an NL2SearchQuery module and a Language-Agnostic RAG (LA-RAG) for document search.
  • A lightweight agentic framework for skill/agent selection, enabling access to documentation, internal toolkits, and code, thereby enhancing chat quality.
  • A robust evaluation and monitoring system to drive continuous improvement and streamline error triaging.
The remainder of the paper is organized as follows: Section 2 covers background information and data sources. Section 3 outlines the overall system design. Section 4 describes the data preprocessing pipelines for knowledge base preparation. Section 5 details the backend infrastructure, and Section 6 explains the advanced document retrieval algorithms. Section 7 presents the frontend components. Section 8 describes the evaluation framework and experimental results. Section 9 discusses related work, and Section 10 concludes the paper.
@@ 2. Background
The primary objective of DECO is to provide a framework for setting up copilots that offer engineers a time-efficient, interactive experience for managing daily tasks and on-call responsibilities. DECO effectively bridges gaps between product telemetry, documentation, on-call duties, code development resources, and team-specific toolkits.
@@ 2.1. Data Sources
By default, DECO incorporates resources from (1) internal documentation repositories, such as troubleshooting guides; (2) historical incident databases; (3) code repositories; and (4) internal Stack Overflow. Teams can also add additional skills to integrate more sources as needed.
@@
Incident data, including information such as title, summary, source, and owning team, is stored in a dedicated database (Chen et al., 2020). This database also captures all communications between engineers, including investigation notes. These records serve not only as a comprehensive documentation resource for incident triaging but also provide invaluable guidance for broader engineering practices.
@@
Similar to IcM databases, Stack Overflow serves as an invaluable resource for Microsoft engineers, offering a wealth of community-driven knowledge and problem-solving techniques. Discussions from these forums are extracted and summarized into reports for future reference. This section provides a comprehensive overview of DECO’s architecture and design decisions.
Figure 2. DECO architecture
@@ 3. Architecture
DECO comprises 4 primary components: Data Preprocessing, Backend Services, Frontend Services, and Evaluation (see Figure 2):
Data Preprocessing: The data preprocessing pipelines are orchestrated in the offline process, executed at user-defined intervals to account for information updates frequently. They draw inputs from various data sources, such as Git, Wikis, and Kusto databases, to generate processed documents stored in the cloud, such as Azure Blob Storage, for online retrieval. Multiple document search indexing services (such as Azure AI Search (Microsoft, 2024b)) are deployed to index the processed documents, enabling seamless access through API calls from downstream backend (online) modules.
Backend Services: The backend orchestrates the agentic framework using PromptFlow (Microsoft, 2024k) for the online interaction with users. It enables straightforward deployment, API and flow monitoring, and greatly reduces maintenance complexity for developers. The workflow can be easily deployed in the cloud, serving as a stateless REST API through managed online endpoint (Microsoft, 2024d) in Azure Machine Learning (Microsoft, 2024c), optimized for real-time, high-volume, low-latency inference requests, and equipped with extensive logging and debugging capabilities, and can be easily scaled to support hundreds of users. Sections 5 and 6 provide further details on these modules.
Frontend Services: The frontend manages user authentication and authorization and handles user interactions. It manages session memory, building the full chat history and sending it to the backend’s (stateless) REST API to enable a continuous conversation experience. Session state and extracted memory data are stored in the cloud, while Azure Log Analytics automatically stores telemetry data for easy monitoring and review from the API calls. This data is later used for feedback learning (Anonymous, 2024b) and evaluation (see Section 8). A web application is deployed with two main user interfaces: a web-based platform and a Microsoft Teams integration, including group chat and session-sharing functionalities.
Compared to many existing agentic frameworks where session and memory management are handled within the same backend (e.g., AutoGen (Wu et al., 2023) or Letta (Packer et al., 2023)), implementing a stateless REST API in the backend to operate independently of session state management and authentication from the frontend, offering two key advantages in production:
@@
Evaluation: A comprehensive assessment of the entire system is conducted to monitor and evaluate real user interactions as well as individual modules, ensuring a high-quality chat experience. The evaluation pipeline supports: (1) Pull Request checks and daily monitoring to track response quality and prevent regressions; (2) leveraging user feedback to construct golden datasets that guide module development and improve response quality; and (3) using user feedback to enhance documentation retrieval, such as through adaptive re-ranking (Anonymous, 2024b).
@@ 4. Data Preprocessing
In this section, we describe the offline preprocessing pipeline for three data sources: IcM (historical mitigated incidents), documentation, and code repositories. These pipelines build the knowledge base (index) to support context-retrieval skills in the backend, such as get_tsg, get_icm, and get_code. Additional pipelines can be created using the same infrastructure to integrate information from other sources, such as Stack Overflow or internal email exchanges.
4.1. Common Infrastructure
DECO provides a unified data preprocessing infrastructure that standardizes ingestion, transformation, and indexing across all pipelines. This common infrastructure includes the following modules, each of which can be customized upon instantiation:
  • Data Ingestion: We use secure authentication to access various sources, such as the IcM database, Git repositories, and documentation sites.
  • Data Transformation: We develop asynchronous scheduled pipelines that extract, and transform raw data to generate structured content in the cloud.
  • Metadata Enrichment: We compute of additional metadata fields, such as reference links and key attributes that can be used to enhance retrieval and re-ranking.
  • Embedding Generation: We use LLMs to generate embeddings across various fields in the structured content.
  • Data Indexing: We create of search indexes based on processed data, with deployment to provide a REST API for efficient querying.